### v1

@@ 4.1 Bend
Bending is a prevalent deformation thanks to the slender morphology of continuum and soft robots (Fig. 5). Herein, we review the evolution of rod models undergoing bending through various mechanisms. Table 1 distills the deformation class reporting representative works.
4.1.1 Bend by Cables
The literature presents several solutions to predict the bending of soft robots. The CRT computed in real-time the kinematics of tendon-actuated robots subject to external loads , modeling the tendon actuation as a single point moment applied to the backbone where each tendon is attached, achieving position errors below on a Nitinol rod (Jones et al., 2009). Direct and inverse kinematic models were presented drawing inspiration from the structure of octopus arms (Giorelli et al., 2012). The direct kinematics exploited a Cosserat rod model. Conversely, the inverse kinematics computes the Jacobian matrix iteratively from the task space. The kinematic models overcome the slack problem as they work directly on the cable tensions instead of controlling the cable extension. Actuations validated the model on a prototype immersed in water. The challenge of predicting the bending angle of a planar manipulator driven by four wires and with triangular notches was addressed using a kinematic Timoshenko rod model (Wenlong et al., 2013). A two-node Timoshenko rod element modeled each triangular unit, accurately predicting the bending angles measured on the tensioned manipulator. Concerning planar grasping, it was proposed a dynamic Cosserat rod model for a tendon-driven continuum robotic finger (Dehghani and Moosavian, 2014b). The finger grasping is simulated as an external point force acting on the finger backbone , assuming that the objects do not slip. Also, a singularity-free version of the model equations is derived using the Taylor expansion. A preliminary experimental validation is conducted by attaching different weights on the tip of a superelastic NiTinol rod. Also, the EBRT was derived using the Lagrangian method to predict the shape of a multi-backbone continuum robot (He et al., 2013). The experimental validation obtained position errors lower than 4%L. Moreover, Gao et al. (2016) predicted manipulator deformations due to external forces, tendon forces, or friction. They provide a planar Cosserat rod model of dexterous continuum manipulators, achieving sub-millimeter tip error predictions in the experimental validation. A dynamic model of a cable-driven soft tentacle was developed to explore the relationship between the strain field and the rope length by formulating a rope-driven actuator model (Ma et al., 2021). A kinematic model based on CRT was presented. Then, a Newton-Euler inverse dynamics algorithm for an equivalent CSM computed the matrices of the Lagrangian inverse dynamics model. A computationally efficient analytical tip force estimator was proposed for tendon-driven catheters (Hooshiar et al., 2021). First, they used a Bezier-based shape interpolation assuming constant curvature. Then, they proposed an optimal problem to invert the system statics and compute the tip force . The experimental validation achieved a N mean absolute error at high frequency. Concerning the interaction with objects, a general CRT-based framework investigated the coupled dynamics of tendon-driven cooperative continuum robots interacting with flexible objects (Jalali and Janabi-Sharifi, 2021). The framework is validated experimentally using two cable-driven spring steel arms performing trajectories with a Nitinol object.
Workspace investigations are crucial for dexterous manipulation tasks. A study on estimating the workspace of CSM was conducted by developing a quasi-static Cosserat rod model and introducing a continuation algorithm for a tendon-driven CSM (Amehri et al., 2021). The algorithm maps the exterior and interior workspace boundaries, addressing challenges like actuator saturation, physical limitations, self-collision, or elastic instabilities. The study highlights that complex configurations of CSM can lead to isolated boundaries and voids. Validation through numerical simulations on planar and spatial configurations demonstrated the effectiveness of the proposal.
Other works focus on the stiffness analysis. A Cosserat rod model for a continuum manipulator with variable stiffness was used to compute the tip stiffness (Zhao et al., 2020a). Experimental results demonstrate the ability to adjust the tip stiffness in various directions with an increase of about ten times the minimum stiffness. Also, a novel cable-driven continuum robot with layer jamming sheaths, which control the stiffness in the transverse and axial directions, is designed, modeled, and validated (Fan et al., 2022). An Euler-Bernoulli rod and the Maxwell-Mohr method derived the effective robot stiffnesses, expressing the rod deflection as
where is the moment due to the unit load. Extensive static force-deflection tests are conducted with different layer jamming sheaths, vacuum pressures, and bending angles. A novel approach proposed additional cables to selectively stiffen planar cable-driven robots with limited coupling with the actuating cables (Molaei et al., 2022). This approach increases the control of shape and stiffness of continuum manipulators without significantly affecting the basic kinematics and addressing uncontrolled deformations like buckling. A planar Cosserat rod model based on (Rucker and Webster III, 2011) identified the design conditions to route the additional cables. The simulations show a contained coupling between the length changes of the actuating and stiffening cables. The proposal is validated experimentally on a prototype with two bending cables and two stiffening cables, demonstrating that the secondary tendons increase the stiffness by ten times. Stiffness is also investigated for a simple rod-shaped silicone soft robot modeled with a geometrically exact CRT (Grube and Seifried, 2022). The stiffness properties are investigated by applying horizontal pulling forces up to 3 N to the rod and robot free-ends. They analyzed experimentally a linear viscoelastic material model and the nonlinear Saint Venant-Kirchhoff material model. The analysis of the damping mechanism has shown that linear damping can be assumed for oscillations with small to medium amplitudes since nonlinear material laws improve the model accuracy only slightly. Therefore, nonlinear effects have to be considered for oscillations of higher amplitudes. Xiao et al. (2023) present an on-demand stiffness modulation approach that retains the intrinsic compliance of elastomeric bodies. They use a Nitinol tube as the backbone, which slides through a cable-driven soft body to achieve robot pose or stiffness modulation. They develop and validate a CRT model in different scenarios by varying and the contact force . Experimental results indicate average shape and tip errors below 1%L, while simulations demonstrate that backbone insertion enhances the workspace and reduces compliance by 57.7%. Their design was further validated by object manipulation and minimally invasive surgery (MIS) case studies.
A generalized kinetostatics CRT model with friction considered for multi-module and multi-segment continuum robots is proposed (Wang et al., 2023). The model is applied to a two-module rod-driven continuum robot with winding ropes to analyze its deformation and load characteristics. The experimental validation, with and without loads, achieves less than shape error.
4.1.2 Bend in Soft Pneumatic Actuators
Several works estimate the bending deflection in pneumatic soft robots. An Euler-Bernoulli rod described a bending pneumatic soft actuator (de Payrebrune and O’Reilly, 2016). Through experimental procedures with the robot clamped at one end, the curvature was expressed as a function of the applied pressure . The results show that the curvature field is a function of the applied pressure.
FEM simulations often provide a benchmark. Using FEM and experiments, a five-parameter constitutive relation for a rod-based pneumatic soft arm model was explored (De Payrebrune and O’Reilly, 2017). The constitutive law’s parameters are determined for pneumatic soft arms through FEM simulations under different conditions. Furthermore, various actuator sections are investigated (e.g., rectangular, semicircular). The parameterization is validated by comparing FEM with the rod model for various actuator sections, subject to tip loads . Another work introduced interoperable dynamic modeling for fluidic elastomeric soft fingers (Mbakop et al., 2021b). The model employs an Euler-Bernoulli framework with Pythagorean Hodograph (PH) curves reconstructing the kinematic shape from virtual control points. The shape of the soft finger is recovered as a function of the applied pressure and control points as
Here, is the height from the ventral part of the finger in the cross-section, and is the normalized curvilinear coordinate along the neutral axis of the soft finger. In this way, a relationship is established between control inputs and PH curve control points, enabling dynamic inversion. FEM simulations validated the PH-EBRT approach for single and two-segment soft fingers, while the experimental validation measures cartesian error and bending curvature on 3D-printed soft fingers of varying stiffness. Interestingly, EBRT successfully modeled a compliant robotic fish tail with fluidic actuators (Wu and Shi, 2022b). The model predicts actuator deflections and nonlinear strains. Nonlinear FEM validated the approach by comparing the results in terms of the first, second, and third obtained natural frequencies, showing good agreement. Finally, the bending of a beam-shaped soft actuator with an embedded fluidic network was analyzed by leveraging a dynamic Euler-Bernoulli rod with large deflections and nonlinear strains (Janizadeh Haji and Bamdad, 2023). Its accuracy is compared against nonlinear FEM and 4th Runge–Kutta. Examining multi-mode and nonlinear strain effects revealed instabilities in some frequencies for the nonlinear model. Geometric parameters influence actuator deformations, with increased channel distribution density enhancing deflection. Moreover, larger fluid channel diameter and actuator length require fewer inlet pressures for the same bending, while higher actuator width and height demand higher pressures. Conversely, Sadati et al. (2019) introduced a ROM and a discretized model based on EBRT with absolute states for the STIFF-FLOP. In addition, they improve the model in (Renda et al., 2018a) with an Euler-Bernoulli discretized model with relative states. An experimental comparison revealed that the ROM is more accurate and numerically robust, while absolute states provide higher computational efficiency. In contrast, relative states show higher sensitivity and computational burden. Finally, a lumped-mass model provided a benchmark with the highest computational cost among the considered models.
Bending structures also exhibit stiffness modulation and adaptation. For instance, CRT modeled a soft bending actuator with an added fabric layer, which changes the actuator stiffness, enhancing the bending and preventing longitudinal stretching (Nalkenani et al., 2021). The authors incorporate fabric effects as a boundary condition in the Cosserat model and validate experimentally across actuation frequencies and loading conditions, producing accurate behavior predictions. Also, it was proposed a nonlinear kinematic Cosserat rod model of a pneumatic surgical CSM with stiffness adaptation (Roshanfar et al., 2022). The model incorporates nonlinear material properties through the two-term Mooney-Rivlin hyperelastic constitutive law, whose strain energy density is
Here, and are the deviatoric strains invariants while and are material constants. The simulation and experimental results for tip displacement and stiffness show similar trends with a maximum error of 8.25%.
More recently, the work in (Namdar Ghalati et al., 2022) presented a method to obtain the static shape of a semicylindrical soft fiber-reinforced pneumatic bending actuator under external force constraints. The pressure-angle relationship in free motion is derived using the Neo-Hookean strain energy theory of hyper-elastic materials, i.e.,
Here, is the first invariant of the right Cauchy-Green deformation tensor and , , are the principal stretches (Odgen et al., 2013). They investigated the effects of external forces using the EBRT, which modeled the arm as a cantilever beam under external load.
4.1.3 Bending with Smart Actuators
Bending is achieved by several innovative actuation technologies. Inspired by the forward crawling motion of the Manduca sexta caterpillar, a dynamic Euler-Bernoulli rod model for a soft robot with shape memory alloy (SMA) actuators was introduced to analyze undulation mechanics (Daily-Diamond et al., 2017). The model employs a motion pattern based on the Witch of Agnesi function
Its shape is analogous to the probability density function of the Cauchy distribution. Lastly, experimental data on an SMA actuator’s flexural rigidity, curvature, and recorded caterpillar leg velocity enable studying ground contact forces and the energetics of successful forward locomotion. Differently, CRT solved direct and inverse kinematics of soft robots actuated by twisted and coiled polymers (Pawlowski and Zhao, 2018). Direct kinematics determines the manipulator shape based on TCP temperatures, and inverse kinematics uses an augmented Jacobian approach (Rucker and Webster, 2011). A subsequent work integrates a thermodynamic TCP model and a dynamic Cosserat rod model, with promising forward dynamics results (Pawlowski et al., 2018). Additionally, an Euler-Bernoulli rod model predicted the deflection of a micro soft robot made from a magnetic elastic composite material (Xiang et al., 2019). The 3D-printed robot, having a head and a tail for rolling and crawling, is controlled by two magnetic fields generated by a 3D Helmholtz coil. Experimental validation for rolling and crawling velocity under different input waveforms, amplitudes, and frequencies shows a linear relation of rolling speed to rotating magnetic field frequency. Moreover, crawling speed correlates positively with magnetic field amplitude and frequency, being faster with sinusoidal compared to triangular wave inputs. A recent work developed a dynamic model of a continuum robotic arm with a 3D-printed elastic backbone and orthogonal plates supporting distributed soft cubic actuators, composed of a novel temperature-responsive hydrogel, each with an embedded Joule heater (Doroudchi et al., 2020). Leveraging the CRT, they simplify the PDE into reduced ordinary differential equations by integrating semi-discretized distributed force and rotational inertia components. Assuming robot isotropy and homogeneity, they estimate the Young modulus and damping coefficients through vibrational tests. Experimental verification accurately predicts tip displacements with a maximum of error. Finally, a study proposes a model of CSM with magnetic actuation (Tariverdi et al., 2020). They use Lie group synchronous variational integrators to derive a geometrically exact model based on an extended dynamic Kirchhoff rod and Hamiltonian formulation. The model considers nonlinearities, magnetic field loads, gravity, and internal and external dissipation forces from friction. Experiments involve dynamic motions with metallic and polymeric soft rods. Polymer rods show higher errors due to fabrication uncertainties and nonlinear elastic properties. Increasing the number of nodes improves the system frequency accuracy.
4.1.4 Bending in Parallel Manipulators
A novel parallel continuum manipulator (PCM) constructed with soft panels and its shape modeling approach was presented (Li et al., 2018). They used the nonlinear EBRT to model the deformations of each soft panel. By solving a constrained optimization for the kinematics, they calculated the deformed shapes and obtained the relationship between the panels’ length and the pose of the moving platform. Experiments validated the model reaching good shape predictions.
Li et al. (2023) present a kinematic model of a planar PCM made of two independent flexible panels. The formulation unifies the geometric constraints, the load balance of the moving platform, and the panel’s statics. For each panel, the relation between internal moment and deflection is described by the EBRT via elliptic integrals, which account for inflection points with large deformations. Simulation results compared against FEM reveal that, under large deflections, the panels exhibit two inflection points close to both ends, which affect the distribution of the internal moment along the panel, (25). Stress checking confirms that the deformations are within the elastic region.
4.1.5 Bending in Tubular Robots
Bending is the prominent deformation of tubular robots aimed for MIS. Pattanshetti and Ryu (2019) employ the CRT to model a laser-patterned monolithic tube with varying structural properties along the length. The combination of CRT and a vision-based identification algorithm enables the characterization of the non-linearities and the bending stiffness along the rod with increasing strain. Another work adjusts the piecewise variable-strain (Renda et al., 2020; Boyer et al., 2020) to simulate multi-section concentric tube robots by including the tubes sliding motion (Renda et al., 2021). Rotation motions of the tubes are included as generalized coordinates instead of boundary kinematic conditions. This procedure led to a minimum set of closed-form algebraic equations solvable for both the shape variables, the actuation forces, and the torques. The resulting equations facilitate control, design optimization, and stability assessment. The work tackles one of the main limitations of the sliding structure of standard CTR, the spaghetti problem.
4.1.6 Bending in Simulated Locomotion
Bending is also exploited to investigate in simulation locomotion strategies (Fig. 5). Zhou et al. (2015b) describe a soft robotic limb attached to a mass with an Euler-Bernoulli rod to examine in quasi-static conditions the peristaltic locomotion controlling the intrinsic curvature on a rough horizontal surface. The paper also provides design recommendations for the optimal performance of soft robots with variable curvature and adhesion. Interestingly, they also studied grasping strategies using a gripper with two fingers attached to a palm modeled using two rods (Zhou et al., 2015a). Concerning snake-like locomotion, especially undulatory swimming, is examined by studying a planar inextensible elastic Cosserat rod with adjustable spontaneous curvature, providing an internal actuation mechanism resembling muscular action in a snake (Cicconofri and DeSimone, 2015). Equations of motion are derived for (a) the rod moving along a prescribed curve and (b) the rod sliding longitudinally without fixed lateral slipping. Non-standard boundary conditions arising from constraints enable solving the equations. The one-dimensional equation captures system kinematics and dynamics without restrictive assumptions on trajectory or actuation, emphasizing the role of spontaneous curvature in driving and steering locomotion.
4.1.7 Bend Miscellaneous
Rods might exhibit interesting and mutable behaviors. Armanini et al. (2017) studied a soft robotic arm clamped at its base with a revolute joint, discovering that when the load is smaller than the buckling critical load, the rod behaves as an elastica compass. The closed curve can be computed analytically by the EBRT. Conversely, when the load exceeds the critical buckling value, the rod behaves like an elastica catapult, causing the snap phenomenon, where the rod releases very quickly elastic potential energy. They computed the dynamic motion using a simplified massless rod model and FEM and validated the behaviors by experimenting with an elastic rod.
Instead, Cacace et al. (2020) present a model for a simulated octopus tentacle based on the dynamics of an inextensible string with curvature constraints and curvature controls. They derive the EoM and appropriate boundary conditions characterizing the corresponding equilibria. The model gives fourth-order evolutive nonlinear controlled PDE, generalizing the classic Euler’s dynamic Elastica equation. They explore the reachability of the model and introduce and compare two optimal control algorithms based on the augmented Lagrangian method solved with adjoint-based gradient descent.
Subclass | References | Rod Model | Contribution |
Cable | Jones et al. (2009) | CRT | Static model for real-time shape estimation of CSM. |
Dehghani and Moosavian (2014b) | CRT | Dynamic model of a continuum finger for planar grasping. | |
Gao et al. (2016) | CRT | Model of notched continuum manipulator for MIS. | |
Pneumatic | De Payrebrune and O’Reilly (2017) | EBRT | Derivation of constitutive law of a planar bending actuator. |
Sadati et al. (2019) | EBRT | ROM and rod models with absolute and relative states validated on STIFF-FLOP. | |
Smart | Daily-Diamond et al. (2017) | EBRT | Dynamic model inspired by caterpillar motion to analyse ondulation mechanics. |
Pawlowski and Zhao (2018) | CRT | Modeling the coupling between CSM body and thermal actuation. | |
Parallel | Li et al. (2018) | EBRT | Shape modeling of a PCM made of soft panels. |
Tubular | Pattanshetti and Ryu (2019) | CRT | Estimation of stiffness variation on a monolithic tube for MIS. |

### v2

@@ 6.1 Bend
Bending is a prevalent deformation in continuum and soft robots due to the natural propensity of these slender structures to bend (Figure 4). Various rod models were employed to predict the bending motion under different actuation mechanisms, stiffness modulation strategies, or dynamic interactions with the environment. To clarify the different modeling approaches, we distinguish between planar (2D) bending and spatial (3D) bending.
6.1.1 Planar bending (2D)
Planar bending assumes that deformation occurs in a single plane, which significantly simplifies the mathematical formulation. Among the approaches, the EBRT is particularly suitable for modeling simple soft robots that bend due to its simplicity. Indeed, EBRT modeled various robotic architectures, achieving sub-millimeter tip position accuracy in multi-backbone continuum robots (he2013analytic) and enabling shape modeling in parallel continuum manipulators via its non-linear formulation (li2018shape). The Timoshenko rod theory, which is suitable for planar deformations (Figure 2), was used to model planar bending in cable-driven soft robots with complex geometries, such as triangular notched structures (wenlong2013mechanics). Pneumatically actuated soft robots also benefited from rod models. Among these, EBRT effectively captured the relationship between internal pressure and bending deformation (de2016constitutive). Later, this approach was extended to handle large deflections and non-linear strains (janizadeh2023steady; wu2022design). sadati2019reduced introduced a ROM and a discretized formulation of the STIFF-FLOP robot using EBRT, demonstrating high accuracy, numerical robustness, and near real-time performances. Further extending EBRT applications, li2023analytic developed a kinematic model for a planar PCM with two independent flexible panels. This formulation integrates geometric constraints, load balance considerations for the moving platform, and panel statics. Beyond kinematic descriptions, gao2016mechanical developed a planar Cosserat rod model for dexterous continuum manipulators, incorporating external forces, tendon forces, and friction, and achieving sub-millimeter tip error predictions in experimental validation. Dynamic planar bending has also been addressed through variational approaches. In particular, tariverdi2020dynamic proposed a Lie group variational integration framework experimentally validated on magnetically actuated soft continuum manipulators, achieving sub-millimeter tip accuracy.
Planar bending models were also applied to investigate locomotion strategies. For instance, zhou2015flexing modeled peristaltic locomotion with an Euler–Bernoulli rod, and cicconofri2015study used a planar Cosserat rod with adjustable curvature to reproduce snake-like swimming. As a further extension, daily2017dynamical developed a dynamic Euler–Bernoulli model of a caterpillar-inspired robot with shape memory alloy (SMA) actuators, linking undulation mechanics to ground contact forces. Similarly, xiang2019study applied an Euler–Bernoulli formulation to predict rolling and crawling in a magnetically actuated soft robot. These works prove the versatility of rod models beyond manipulation.
6.1.2 Spatial bending (3D)
Spatial bending occurs when deformation is not confined to a single plane but involves curvature around multiple axes, eventually accompanied by torsion. These cases typically require Cosserat or Timoshenko formulations to accurately capture the complete range of strain modes. CRT, despite its higher complexity, has also extensively contributed to bending modeling in cable-driven robots (jones2009three_dim). An early example of CRT-based bending modeling is provided by giorelli2012two, who developed a kinematic model for a cable-driven soft arm with non-uniform cross-sections and derived its inverse model using the Jacobian matrix.
In pneumatic actuators, CRT was coupled with nonlinear constitutive laws (e.g., Mooney–Rivlin hyperelasticity) to improve accuracy (Roshanfar2023Hyperelastic). Similarly, namdar2022static computed the static shape of a semicylindrical soft fiber-reinforced pneumatic bending actuator under external force constraints, using a Neo-Hookean model to derive the pressure-angle relationship in free motion (ogden2013). CRT was also applied to soft robots embedding smart actuators (Pawlowski2018modeling; Pawlowski2018dynam_model; doroudchi2020dynamic). Beyond kinematic descriptions, dynamic models mainly based on CRT, provide more suitability for dynamic tasks and allow the integration of actuation into its formulation (Ma2021dynamicsModel). Further refinements in CRT-based cable-driven robot models included friction effects arising from interactions between cables and the robot structure (wang2023general). CRT was also leveraged to model tasks involving interaction with external objects. jalali2021dynamic introduced a CRT-based framework to analyze the coupled dynamics of tendon-driven cooperative continuum robots interacting with flexible objects. Likewise, dehghani2014dynamics proposed a dynamic Cosserat rod model for a tendon-driven continuum robotic finger, where the grasping forces are modeled as external forces acting on the backbone. These studies highlight the effectiveness of CRT in dynamic interaction tasks, including force estimation in cable-driven catheters, as demonstrated by hooshiar2021analytical, who used a Bézier-based shape interpolation and an optimization approach to estimate tip forces with high precision. Another important aspect of dynamic interactions concerns workspace estimation, namely the determination of the spatial region that the soft robot’s end-effector is able to reach given its actuation and mechanical constraints. amehri2021workspace developed a quasi-static CRT model with a continuation algorithm to map workspace boundaries in tendon-driven manipulators, addressing self-collision, elastic instabilities, and actuator saturation. The study found that complex CSM configurations could lead to isolated boundaries and voids, which are essential to consider for accurate workspace predictions. Beyond kinematic and dynamic predictions, stiffness modulation also plays an important role in predicting soft robots behavior (xiao2023kinematics) as well as mitigating undesired effects such as buckling, which can hinder precise movements (molaei2022cable). Notable advancements involve variable stiffness continuum manipulators, which, leveraging CRT-based modeling can increase stiffness up to ten times, resulting in improved control precision (zhao2020sim). Other techniques include the addition of layer jamming sheaths to control stiffness both along the axial and transverse directions (fan2022novel). In parallel, research on a simple rod-shaped silicone soft robot, modeled using CRT, investigated stiffness properties by applying horizontal pulling forces, up to 3 N, at the free ends of the rod (grube2022simulation). Two material models were examined: a linear viscoelastic model and a nonlinear Saint Venant-Kirchhoff material model. The damping mechanism was analyzed, revealing that linear damping can be assumed for small to medium oscillations. However, non-linear effects become relevant when oscillations reach higher amplitudes, enhancing the model’s accuracy. Nalkenani2021modeling presented a CRT model of a soft bending actuator with an additional fabric layer to modify the actuator stiffness, modeled as boundary condition of the Cosserat model. This adjustment improved bending behavior while preventing longitudinal stretching.
FEM was widely used to augment and validate rod-based models, particularly for pneumatic actuators. For instance, de2017development proposed a five-parameter constitutive relation for a pneumatic soft arm, with parameters identified via FEM simulations under different loading conditions. Conversely, mbakop2021inverse integrated Pythagorean Hodograph (PH) curves with an Euler–Bernoulli framework to reconstruct actuator shapes from virtual control points, establishing a direct link between control inputs and geometry. Both approaches, as well as more recent formulations (li2023analytic), were validated through FEM and experimental results, confirming their accuracy. In parallel, rone2013continuum developed a high-fidelity lumped-parameter dynamic model for cable-driven continuum robots, based on the principle of virtual power. It captured curvature variations while accounting for inertial, actuation, friction, elastic, and gravitational effects.